Avoiding a collision. A car is traveling on a road that is perpendicular to a railroad track. When the car is 30 meters from the crossing, the car's new collision detector warns the driver that there is a train 50 meters from the car and heading toward the same crossing. How far is the train from the crossing?
Answer by stanbon(48568)

A car is traveling on a road that is perpendicular to a railroad track. When the car is 30 meters from the crossing, the car's new collision detector warns the driver that there is a train 50 meters from the car and heading toward the same crossing. How far is the train from the crossing?
Draw the picture.
It is a right triangle with a side of 30 and hypotenuse of 50.
Use Pythagoras to find the other side.
50^2=30^2+x^2
x^2=2500-900=1600
x=400 meters (the distance the train is from the crossing)
Cheers,
stan H.